Using stokes theorem (No point otherwise) find the F.dt, F vector=<z,-z,x^2-y^2> and C is the three lines in which z=8-4x-2y (plane) cuts the coordinate planes.5. USING STOKES THEOREM (NO POINTS OTHERWISE, FIND FocF, F={z LINES IN WHICH 258-4xby CIMKE COORDINATE PLANES (2, -2, x-x) AN

Please be detail, thanks.
Answer #1

using stockes theorem find &ridi 14 zł -23 + x2-422 $7 C: is the theree lines in which cuts the co-ordinate Z-08-42-24 By usiNow let Чz +29 +2 8 - 4 + 23 + үф Пуф) р + 4 1+23 16 + 4 + р 4T + 2 + 2 эн How we ад projection on x-y plane- Z – Чx + 29 ә2.projection on ty plane Z=0 Y 4x+2y=0 22 + y = 4 424 2x+y=4 Now (Gx . ñ ds SS (SXF) Sf4-4+1) – -1) 1) 47 +23+ K x Jav dedy - T$? (4-2x) C4x10) dx 2 (-8x.3 + 362 – 40) dx 03 -542 +1207 3 86 - 08 3 29.33 Hena ง 3 Sf(xXF).Ônds 3 By stockes theoreu $7.07

If any doubt ask in comment box..thank you....